please answer all questions Beers Law Plot for KMnO, solution 0.8 0.7 06 05 0.4 03 02 0.1 y -4.84E 03x-2.26E-01 R 997E 01 KMnO4 Concentration 6.40× 10-5 9.60× 10-5 1.28 x 104 0.383 1.60× 10- 1.92 10- 0.093 0.228 0.570 0.697 0 5 00E 05 800E 05 1.10E 04 1.40E 04 1.70E 04 2.00E 04 concentration, M 1. The calibration curve and the calibration equation above was obtained by ploting absorbance plotting absorbance s. [KMnO4]. In the calibration equation y 4.84E+03 -2.26E-01, what does y and x represent, respectively? 2. If the absorbance of a KMn0, solution of unknown concentration is 0.833, calculate the concentration of KMnO4 in the solution. Show your calculation.
Answer #1

1.

y is absorbance

x is molar concentration of KMnO_4

2.

\\ \: \\ \: y = 4.84E+03x-2.26E-01 \\ \: \\ \: 0.833= 4.84 \times 10^{+03}x-2.26 \times 10^{-1} \\ \: \\ \:1.059 = 4.84 \times 10^{+03}x \\ \: \\ \:x =0.000219 \: M

The molar concentration of KMnO_4 is 0.000219 M.